3 Catching Lobsters Suppose the best lobsters are sold on the market for 50 However, the overall quality of a catch drops as more lobsters are caught. So, the total benefit to catching a pound lobsters is TB 50Q- 2 and the marginal benefit is MB-50-Q. The marginal cost of catching a pound of lobsters is 30 9. (1 point) Suppose the lobster fishery is privately-owned. How many lobsters will the owner catch? 10. (1 point) Suppose the lobster fishery is open access. How many lobsters will be caught by all the fisherfolk?
Answer #1

9. Private firm will catch till point where MB = MC.
50 - Q= 30
Q = 20

10. In open access, fishing is done it AB = MC
AB = 50 - 1/2Q = 30
Q = 40